OpenStudy (anonymous):

Suppose that u and v are differentiable at functions of x, use the given values of the functions and their derivitaves to find the value of the indicated deriv. u(1)= 5, u'(1)=-5, v(1)=6, v'(1)=-2

OpenStudy (anonymous):

the derivative of a function f(u,v) is equal to d/du f(u,v) * u'(t)+ d/dv f(u,v) * v'(t) at time t=1, this is equal to d/du f(5,6) * (-5) + d/dv f(5,6) * (-2)

OpenStudy (anonymous):

if f(u,v)=u*v, then d/du f(u,v)=v and d/dv f(u,v)=u 6*(-5)+5*(-2)

OpenStudy (anonymous):

which is just the product rule